A foreman for an injection-molding firm admits that on 10% of his shifts, he forgets to shut off
the injection machine on his line. This causes the machine to overheat, increasing the probability
that a defective molding will be produced during the early morning run from 2% to 23%. If a
molding is randomly selected from the early morning run of a random day, what is the probability
that it is defective?

Answer by ikleyn(52752)

.

In  10%  of cases the  23%  rule works;


in the rest  90%  of cases the rule of 2%  works.



THEREFORE, the total average is  0.2*0.23 + 0.9*0.02 = 0.046 + 0.018 = 0.064 = 6.4%.    ANSWER

Solved, answered and explained.